OpenStudy (vane11):

I forgot how to do this it's like algebra level so kinda embarrassing: Find the slope intercept form of a line satisfying the given conditions; passing through (-2,4) and parallel to a line passing (-5,2) and (-3,-1)

jimthompson5910 (jim_thompson5910):

first you need the slope of the line through (-5,2) and (-3,-1)

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-1-2)/(-3-(-5)) m = (-1-2)/(-3+5) m = ???

OpenStudy (vane11):

-3/2

jimthompson5910 (jim_thompson5910):

so we know the slope is m = -3/2 and we want this line to go through (-2,4)

jimthompson5910 (jim_thompson5910):

y = mx + b y = -3/2x + b 4 = (-3/2)(-2) + b 4 = 6/2 + b 4 = 3 + b ... b = ??

OpenStudy (vane11):

1

jimthompson5910 (jim_thompson5910):

you got it, so the equation is y = -3/2x + 1

OpenStudy (vane11):

oh, wow! haha that was a lot easier than what I thought haha thankyou! And thanks for showing me step by step not just gving me the answer

jimthompson5910 (jim_thompson5910):

you're welcome